OpenStudy (calculusxy):

Medal for best answer! A store has a discount of 15% off of all items. Amy has a $50 gift certificate and she wants to buy a jacket for $50 and a second item. What is the maximum original price of the second item?

OpenStudy (calculusxy):

@perl

OpenStudy (calculusxy):

@amistre64 @jim_thompson5910

OpenStudy (dan815):

I think they mean first item?

OpenStudy (anonymous):

So are both items 50% of or the final purchase?

OpenStudy (anonymous):

*off

OpenStudy (calculusxy):

The discount is only 15% off of all the items in the store.

OpenStudy (calculusxy):

The rest I am not too sure about.

OpenStudy (amistre64):

is all she has to spend is the 50 gift certificate?

OpenStudy (amistre64):

50(1-.15) for the jacket, leaves 50(.15) that can be covered

OpenStudy (calculusxy):

Yes. That's what I think so. It didn't give much more information about it.

OpenStudy (perl):

the jacket* costs .85 * 50 , so how much money does she have left

OpenStudy (calculusxy):

$42.50 and if you subtract that from $50 it is $7.50 ?

OpenStudy (amistre64):

K(1-.15) = 50(.15)

OpenStudy (perl):

.85x = 7.50 x is your maximum price object

OpenStudy (dan815):

ooh sheesh , the question sounds like the jacket is 50 dollars after the 15% discount so it didnt make sense.. why she would even have money for a 2nd item -.-

OpenStudy (amistre64):

she buys 2 things, each one is 15% off, so she pays: A(1-.15) + K(1-.15) = 50 the first item is 50 soo find K 50(1-.15) + K(1-.15) = 50 sounds like a good approach

OpenStudy (perl):

I read it as, both the jacket and the second item are 15% off.

OpenStudy (calculusxy):

But doesn't it say "maximum ORIGINAL amount?"

OpenStudy (calculusxy):

I don't think original can have any discounts in it.

OpenStudy (kainui):

@calculusxy Yeah that's right

OpenStudy (calculusxy):

So here is what I did. \[0.85($50) + x \le $50\] \[$42.50 + x \le $50\] \[x \le $7.50\]

OpenStudy (amistre64):

x is on sale, so your only paying .85x not x

OpenStudy (calculusxy):

x is the second item

OpenStudy (amistre64):

of course it is ....

OpenStudy (calculusxy):

yes so that is what I am trying to find out , after all it says original

OpenStudy (amistre64):

A store has a discount of 15% off of all items. We are buying 2 things: A and K Amy has a $50 gift certificate .85(A+K) = 50 and she wants to buy a jacket for $50 and a second item. .85(50+K) = 50 What is the maximum original price of the second item? ^^^^^^ these are A and K

OpenStudy (calculusxy):

But why would you put 15% on both of the items?

OpenStudy (amistre64):

why would the information say that ALL THINGS in the universe are 15% off?

OpenStudy (calculusxy):

But it is saying the "maximumORIGINAL amount."

OpenStudy (calculusxy):

That is what threw me off.

OpenStudy (amistre64):

yep, its does.

OpenStudy (amistre64):

what is the price of the item seeing that its sale price (not its original price) is what you are paying for.

OpenStudy (calculusxy):

Of course I would've taken 15% off if it didn't say that. It would seem pretty obvious for me to do that. But that part just my nerves and confused me and several other classmates of mine.

OpenStudy (amistre64):

spose we but 3 things .85(a+b+c) for $50 abc are original prices

OpenStudy (calculusxy):

okay...

OpenStudy (amistre64):

spose we want to buy out the whole store .... we would purchase it all of85% of its original value

OpenStudy (amistre64):

does this make sense yet?

OpenStudy (calculusxy):

yeah

OpenStudy (amistre64):

we have 7.50 left to cover an item thats on sale .85x = 7.50 is all we can afford

OpenStudy (calculusxy):

i did not get that part. sorry

OpenStudy (amistre64):

is the item we are buying on sale?

OpenStudy (calculusxy):

yes the items we are buying are on sale.

OpenStudy (amistre64):

then is .85x its sale price?

OpenStudy (calculusxy):

yeah

OpenStudy (amistre64):

then we dont want to know its sale price, we want know its original price: x so solve for x

OpenStudy (calculusxy):

x=6.375?

OpenStudy (amistre64):

no, when something goes on sale it doesnt generally cost MORE to purchace

OpenStudy (amistre64):

.85x = 7.50 divide both sides by ... ,85

OpenStudy (calculusxy):

8.82?

OpenStudy (amistre64):

much better

OpenStudy (amistre64):

so some item x <= 8.82 will be covered by the remaining balance on the gift card

OpenStudy (calculusxy):

okay that makes much more sense. thank you so much!

OpenStudy (amistre64):

youre welcme

OpenStudy (calculusxy):

my friends and I are having some misunderstanding wid this question. thanks for going it over.